katelyn2002:

Zach is 4 years older than twice his sister Maya's age. If Maya's 5 years old how old is Zach?

Eiwoh2:

Simple math. This can be done with a calculator, but if these calculations are made, Zach would be 14 years old. Twice his sister's age, 5, would be 10, and adding 4 more would be 14.

theyhateme:

Z = 2m + 4 = 2(5) + 4 = 14 years old ---> answer